# Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents

### v3

@@ 4.3 Formal-LLM Prompts and Planning from Automaton
The automaton initiates the plan generation process from the start state. In situations where multiple viable options are available for the automaton to proceed, a prompt is created to inquire the LLM, as shown in Figure 1. This prompt encompasses a comprehensive {task description} to mitigate potential forgetting issues of the LLM (Hua et al., 2023), details about the {current progress}, and {feasible choices} determined by the PDA. For example, consider the PDA in Figure 3 when it is in state $q_{1}$ with the symbol $I$ at the top of the stack. There are three viable transits: $(\varepsilon,I;AI)$, $(\varepsilon,I;CT)$, and $(i,I;\varepsilon)$. In practice, as nonterminals $A$ and $C$ are promptly replaced by $terminals$ from $a_{1}\sim a_{4}$ and $c_{1}$, we directly use the tool name (e.g., Colorization) instead of the category name (e.g., “Image-in, Image-out tools”) in the prompt. This is because the tool name provides more concrete information for the LLM to comprehend the tool’s functionality, enabling the LLM to make informed decisions without influencing the plan’s executability. A prompt example is provided in Appendix B.3.
Another advantage of employing PDA for LLM-based agent planning is its ability to articulate plans that involve tools with many-input-single-output (e.g., Image-Text pair as input, Text as output), assuming the number of inputs of each tool is deterministic. To illustrate, we depict the derivation tree of the string $e_{1}a_{1}ib_{1}i$ in Figure 5. Treating the tools $a_{1}$ and $b_{1}$ as functions, the string is functionally equivalent to $t=e_{1}\big{(}a_{1}(i),b_{1}(i)\big{)}$. When the PDA finishes the substitution of symbols $E$ and $I$, it can well know that the stack symbol $T$ is derived from $(\varepsilon,T;EIT)$ and matches the input of tool $e_{1}$, because they are pushed into the stack at the same time. Consequently, we can readily identify the tools that will operate on this image/text data in the subsequent step and incorporate this information in the prompt, enabling the LLM to make more informed decisions.
Figure 3: An equivalent PDA to the CFG of the combination of Eq.(3) and Eq.(4). $\varepsilon$ represents the empty string, and the lowercase letters $a$ to $f$ are $a_{i}$ to $f_{j}$ in Eq.(3). We take the transit $(b,B;\varepsilon)$ as an example. It implies that if the PDA is at state $q_{1}$, the top of the stack is $B$, and the next letter of the input string is $b_{1}$, $b_{2}$ or $b_{3}$, then the PDA can remove the stack top $B$ and remain in state $q_{1}$.
Figure 4: The PDA includes 13 states, $\{10,11,...,19,20,Start,Fin\}$, with $Start$ as the start state and $\{Fin\}$ as the accept state set. The lowercase letter $a$ is for breakfast, $b$ is for lunch, $c$ is for supper, $d$ is for playing basketball, $e$ is for grocery shopping, $f$ is for house cleaning, $g$ is for homework, and $h$ is for laundry. The stack symbols $D,Z,E,N$ represent that there are 3, 2, 1, and 0 meals remaining to plan, respectively. $\forall$ represents any stack symbol. The variable $i$ is enumerated from 20 to 11, $j$ is enumerated from 20 to 11 except 13, and $k$ is enumerated from 20 to 13.
@@
4.4 Backtracking Mechanism and Reinforcement Learning from Task Feedback
@@
Upon imposing constraints on the tool usage, a new challenge emerges: the automaton may encounter a dead-end during plan generation. Consider the formal language in Eq.(3) and Eq.(4). Upon analyzing the language, we observe that the only means of transitioning from the symbol $T$ to $I$ (and subsequently to a terminal $i$) involve $B$-type (Image-in, Text-out) and $E$-type (Image-Text-pair-in, Text-out) tools. If these two tool types are exhausted, yet there are still $T$ symbols in the stack, the automaton will inevitably fail to reach the accept state $q_{2}$, as there are no remaining methods to clear the symbol $T$. Consequently, a dead-end situation may arise following the imposition of tool usage limits.
To address this challenge, we propose to document the automaton’s details at each step, encompassing the current state and stack, and the part of the word that has been generated. When the automaton confronts a dead-end (indicating no viable next step), we initiate backtracking to the preceding steps until there is an unexplored branch. The backtracking involves retracing the automaton to the prior step, reinstating the automaton details based on the recorded stack, and eliminating the dead-end branch from the prior choice list. Consequently, the automaton is capable and guaranteed to generate a valid, executable plan.
The design above only guarantees the validity of the generated plans. For example, the plan $e_{1}a_{1}ib_{1}i$ shown in Figure 5 is an executable plan for the task “Given blurry grayscale images, how to return the object names in English step by step?”. However, this may not be the optimal plan. Thus, as an additional enhancement in plan quality, particularly when employing open-source LLMs, we integrate Reinforcement Learning from Task Feedback (RLTF) (Ge et al., 2023a) following the application of our Formal-LLM framework. After the generation of a plan by the LLM-based agent for a training task, the plan is executed on benchmark data to assess its performance. Then, the performance is used as a reward for Reinforcement Learning (RL) to update the parameters of the LLM. Our framework ensures that invalid plans are always excluded during the LLM-based agent’s plan generation. Therefore, the number of valid rewards for LLM fine-tuning increases, and our approach can improve the performance of fine-tuned LLM-based agents.
@@ 5 Experiments
5.1 Backbone LLMs
We test our Formal-LLM framework on both closed-source and open-source LLMs. Closed-source LLMs include:
  • GPT-3.5-turbo (Brown et al., 2020) is a generative pre-trained transformer of OpenAI with a strong ability of various language tasks based on instructions.
  • Claude-2 (Claude-2, 2023) is a transformer LLM of Anthropic trained with unsupervised learning and Reinforcement Learning from Human Feedback (RLHF).
  • GPT-4 (OpenAI, 2023) is a follow-up version of GPT-3.5, which is more powerful than its predecessors.
Open-source LLMs include:
  • Flan-T5-Large (Chung et al., 2022) is an encoder-decoder language model with 770 million parameters, pre-trained by instruction fine-tuning.
  • Vicuna-7B (Chiang et al., 2023) is a decoder-only chatbot trained by fine-tuning the LLaMA model (Touvron et al., 2023a) with user-shared conversations.
  • LLaMA-2 (Touvron et al., 2023b) is a successor to the original LLaMA model. It is a more powerful language model. In this work, we use the 13-billion size model.
@@ 5.2 Learning Schema of LLMs
  • Few-shot Learning (Few) presents a set of high-quality demonstrations in the prompt, each consisting of both input and desired output on the target task.
  • Reinforcement Learning from Task Feedback (RLTF) executes a plan and takes its performance as the reward to optimize the LLM through an RL process.
  • Formal-LLM (F-LLM) is our proposed framework, leveraging automaton to control the LLM’s generation.
  • Formal-LLM plus RLTF (F-LLM+RLTF) applies our Formal-LLM framework on top of RLTF, which excludes invalid plans and increases the amount of valid rewards for RLTF.
Specifically, we utilize Zero, Few, and F-LLM frameworks for closed-source LLMs, as these frameworks do not need to modify the LLM’s parameters. Regarding open-source LLMs, we compare RLTF, F-LLM and F-LLM+RLTF, given that the RLTF method surpasses Zero and Few, as exhaustively discussed in recent work (Ge et al., 2023a).
5.3 Experimentation Datasets
5.3.1 Benchmark Datasets
For the benchmark experiment, we utilize the tools and multi-step tasks in the OpenAGI platform (Ge et al., 2023a). The tools are detailed in Table 3 in the Appendix, and the tasks are categorized based on their output type and ground-truth label type (Task 1, 2, and 3). Then, based on different task types, different metrics are employed to gauge the performance: CLIP Score (Hessel et al., 2021), assessing the similarity between text and image, is utilized for Text-to-Image tasks; BERT Score (Zhang et al., 2020), evaluating text generation with BERT, is applied when both data labels and the expected outputs are texts; and ViT Score (Wu et al., 2020) gauges the similarity between the image label and image output. Additionally, we construct Task X, which is a subset of “Task 1 $\cup$ Task 2 $\cup$ Task 3” that require tree-structured plan rather than chain-structured plan due to the use of many-input-single-ouput tools, such as Visual Question Answering. Task X is used to test the complex planning ability of our Formal-LLM framework. The data samples in Task X are evaluated using their corresponding metric. Example tasks of each category are illustrated in Table 4 in the Appendix.
5.3.2 Real-life Practical Tasks
We also experiment with real-life planning scenarios, encompassing daily plans, cooking recipes, and risk management, where validity and rationality are crucial. In these scenarios, the concept of tool is generalized to include various types of steps in a plan, such as events, actions, or activities, as they can take diverse forms supporting the execution of the plan. For example, a tool in the daily planning task may be an activity such as taking breakfast. We provide qualitative analyses between the Zero and F-LLM learning schema using GPT-4 backbone for practical tasks. We use GPT-4 for the experiment of practical tasks because other LLMs can hardly generate readable plans for these tasks, and we test under Zero and F-LLM due to the limited sample size and that these two learning frameworks do not require fine-tuning or accessing to the LLM parameters.
@@ 5.4 Experimental Analysis
The experiment results on the benchmark tasks are shown in Table 1 and Table 2, referring to the closed-source and open-source LLMs, respectively. Each row stands for a type of task, each column represents the learning schema of an LLM-based agent, and every three columns are the results of the same LLM. For closed-source LLMs, among the three learning schemata without revising the LLM parameter, almost all the best scores under each type of task belong to our F-LLM framework. For open-source LLMs, our F-LLM without fine-tuning is already better than RLTF (the best schema in the OpenAGI platform (Ge et al., 2023a)) in most cases, except for the ViT Score (Task 3), because the Type-3 tasks take the most significant portion of the RLTF’s fine-tuning data, and thus RLTF is adequately optimized on these tasks. The performance gain of our F-LLM framework comes from the 100% executable plans and supporting tree-structured planning for the difficult tasks in Task X. As a comparison, the best open-source LLM (LLaMA-2-13B with RLTF) can only generate 47% executable plans, and the best closed-source LLM (GPT-4 with few shot) can generate 76% executable plans. Additionally, none of the baselines can handle the Task X (Score = 0.0). Fortunately, due to the increased amounts of valid rewards from the 100% executable plans with our F-LLM framework, the F-LLM+RLTF approach enables boosted performance, as demonstrated by its better scores in Table 2.
To summarize, the benchmark experiment results show that our F-LLM framework is an effective method to integrate the benefits of both natural and formal languages for more controllable and valid planning of LLM-based agents.
@@ 5.5 Case Study
We also show the results of real-life examples by applying our Formal-LLM to the GPT models. Due to the page limit, we put the complete results in the Appendix.
From our observation, after applying our Formal-LLM, the generated plan from the GPT-based agent is more complete, reasonable and specific to the case. In the daily plan example, the agent fails to fit all the activities into the plan without the Formal-LLM, while it can achieve this goal by applying the Formal-LLM. Still take Figure 4 as an example, our Formal-LLM framework can limit the planning to 10:00$\sim$20:00, while without Formal-LLM, the agent schedules activities after 20:00, even if we mention the constraint “Generate a plan for all the activities between 10:00 and 20:00” in natural language, which shows the advantages of formal language guided planning. For the cooking recipe example, without strict automaton constraint, GPT may generate steps such as “stir-fry the Chinese broccoli until it changes color” after “once the water boils, add the Chinese broccoli”. This is unreasonable because it’s impossible to stir-fry when broccoli is in boiling water. As for the risk management example, the generated plan without our framework is too general that it could be used for any other two companies. However, the plan after applying Formal-LLM is more specific and focuses on the potential antitrust risk between Microsoft and Blizzard. Thus, after applying the automaton which describes the constraints, the generated plan has better quality under the Formal-LLM framework.
@@ 6 Conclusions and Future Work
In this study, we introduce the innovative Formal-LLM framework for LLM-based agents, combining LLM’s power of natural language comprehension with the precision of formal language. Our experiments, encompassing benchmark tasks and real-life practical scenarios, affirm the feasibility and effectiveness of employing automaton to control the agent’s generation of valid plans. More controllable LLM-based agents can augment the potential utilization of LLM in applications where high validity of planning holds significant importance.
Several potential extensions to our work merit consideration. First, automating the translation of natural language into formal language for agents could further improve the framework, as designing formal language and automata by domain experts may not be the most efficient approach. Additionally, this work focuses on LLM plan generation based on formal language; however, another important problem to explore in the future is LLM plan verification based on formal language. Furthermore, our framework uses one large automaton to control the whole planning process, while another potential solution is to use multiple small automata, each describing one constraint, and the agent only accesses the corresponding one when needed during plan generation.
ACKNOWLEDGMENTS
We thank Harun Taha Kepenek for helpful suggestions on improving our daily plan task.
Appendix A Impact Statements
11footnotetext: https://github.com/richzhang/colorization22footnotetext: https://huggingface.co/caidas/swin2SR-classical-sr-x2-6433footnotetext: https://github.com/swz30/Restormer44footnotetext: https://huggingface.co/google/vit-base-patch16-22455footnotetext: https://huggingface.co/facebook/detr-resnet-10166footnotetext: https://huggingface.co/nlpconnect/vit-gpt2-image-captioning77footnotetext: https://huggingface.co/CompVis/stable-diffusion-v1-488footnotetext: https://huggingface.co/distilbert-base-uncased-finetuned-sst-2-english99footnotetext: https://huggingface.co/facebook/bart-large-cnn1010footnotetext: https://huggingface.co/t5-base1111footnotetext: https://huggingface.co/distilbert-base-uncased1212footnotetext: https://huggingface.co/gpt21313footnotetext: https://huggingface.co/microsoft/git-base-textvqa1414footnotetext: https://huggingface.co/distilbert-base-cased-distilled-squad
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix B Appendix
B.1 Bechmark Tasks and Tools
@@
Image Image ($A$) Colorization ($a_{1}$) ColorizerA (Zhang et al., 2017)
Super-Resolution ($a_{2}$) Swin2SRA (Conde et al., 2022)
Image Denoising ($a_{3}$) RestormerA (Zamir et al., 2022)
Image Deblurring ($a_{4}$) RestormerA (Zamir et al., 2022)
Text ($B$) Image Classification ($b_{1}$) ViTA (Dosovitskiy et al., 2021)
Object Detection ($b_{2}$) DETRA (Carion et al., 2020)
Image Captioning ($b_{3}$) Vision Encoder DecoderA (Li et al., 2023b)
Text Image ($C$) Text-to-Image Generation ($c_{1}$) Stable DiffusionA (Rombach et al., 2022)
Text ($D$) Sentiment Analysis ($d_{1}$) DistilBERTA (Sanh et al., 2020)
Text Summarization ($d_{2}$) BARTA (Lewis et al., 2019)
Machine Translation ($d_{3}$) T5A (Raffel et al., 2020)
Fill Mask ($d_{4}$) DistilBERTA (Sanh et al., 2020)
Text Generation ($d_{5}$) GPT-2A (Radford et al., 2019)
Image-Text Pair Text ($E$) Visual Question Answering ($e_{1}$) GITA (Wang et al., 2022)
Text-Text Pair Text ($F$) Question Answering ($f_{1}$) DistilBERTA (Sanh et al., 2020)
@@
B.2 Implementation Details
Our framework and all baselines are implemented by PyTorch, an open-source library. We follow the implementation setting of the OpenAGI platform (Ge et al., 2023a) for baselines. For the F-LLM+RLTF framework, we use the REINFORCE (Williams, 1992) as the core reinforcement learning algorithm of RLTF. We use the original checkpoint of each backbone LLM without supervised fine-tuning. We set the maximum number of updating epochs at 30 and use Adam as the optimizer with the learning rate at 0.001 for RLTF. Also, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to the RLTF for efficient fine-tuning.
B.3 Prompt for Benchmark Tasks
An example of the CFG and the corresponding PDA are shown in Eq.(3) and Eq.(4) as well as Figure 3. Different benchmark tasks share a common subset of constraints shown in Eq.(3) corresponding to the input-output constraints of the tools. Besides, each task has it own contraint on the input-outpt of the task, resulting in different contraints in Eq.(4), and thus the final PDA (Figure 3) for different tasks are different.

### v4

@@ 4.3 Formal-LLM Prompts and Planning from Automaton
The automaton initiates the plan generation process from the start state. In situations where multiple viable options are available for the automaton to proceed, a prompt is created to inquire the LLM, as shown in Figure 1. This prompt encompasses a comprehensive {task description} to mitigate potential forgetting issues of the LLM (Hua et al., 2023), details about the {current progress}, and {feasible choices} determined by the PDA. For example, when the PDA in Figure 3 is in state $q_{1}$ with the symbol $I$ at the top of the stack, three viable transits: $(\varepsilon,I;AI)$, $(\varepsilon,I;CT)$, and $(i,I;\varepsilon)$ are included in the {feasible choices} of the prompt. A prompt example is provided in Appendix A.3.
Another advantage of employing PDA for LLM-based agent planning is its ability to articulate plans that involve tools with many-input-single-output (e.g., Image-Text pair as input, Text as output), assuming the number of inputs of each tool is deterministic. To illustrate, we depict the derivation tree of the string $e_{1}a_{1}ib_{1}i$ in Figure 5. Treating the tools $a_{1}$ and $b_{1}$ as functions, the string is functionally equivalent to $t=e_{1}\big{(}a_{1}(i),b_{1}(i)\big{)}$. When the PDA finishes the substitution of symbols $E$ and $I$, it can well know that the stack symbol $T$ is derived from $(\varepsilon,T;EIT)$ and matches the input of tool $e_{1}$, because they are pushed into the stack at the same time. Finally, we can identify the tools that will operate on this image/text data in the subsequent step and incorporate this information in the prompt, enabling the LLM to make informed decisions.
Figure 4: The PDA includes 13 states, $\{10,11,...,19,20,Start,Fin\}$, with $Start$ as the start state and $\{Fin\}$ as the accept state set. The lowercase letter $a$ is for breakfast, $b$ is for lunch, $c$ is for supper, $d$ is for playing basketball, $e$ is for grocery shopping, $f$ is for house cleaning, $g$ is for homework, and $h$ is for laundry. The stack symbols $D,Z,E,N$ represent 3, 2, 1, and 0 meals remaining to plan, respectively. $\forall$ represents any stack symbol. The variable $i$ is enumerated from 20 to 11, $j$ is enumerated from 20 to 11 except 13, and $k$ is enumerated from 20 to 13.
@@
4.4 Reinforcement Learning from Task Feedback and Backtracking Mechanism
@@
Upon imposing constraints on the tool usage, a new challenge emerges: the automaton may encounter a dead-end during plan generation. Consider the formal language in Eq.(3) and Eq.(4). Upon analyzing the language, we observe that the only means of transitioning from the symbol $T$ to $I$ (and subsequently to a terminal $i$) involve $B$-type (Image-in, Text-out) and $E$-type (Image-Text-pair-in, Text-out) tools. If these two tool types are exhausted, yet there are still $T$ symbols in the stack, the automaton will inevitably fail to reach the accept state $q_{2}$, as no methods remain to clear the symbol $T$. Finally, a dead-end situation may arise following the imposition of tool usage limits.
To address this challenge, we propose to document the automaton’s details at each step, encompassing the current state and stack, and the part of the word that has been generated. When the automaton confronts a dead-end, we initiate backtracking to the preceding steps until there is an unexplored branch. The backtracking involves retracing the automaton to the prior step, reinstating the automaton details based on the recorded stack, and eliminating the dead-end branch from the prior choice list. Thus, the automaton is capable and guaranteed to generate a valid plan if it exists.
The design above only guarantees the validity of the generated plans, but not the optimal plan. Thus, as an additional enhancement in plan quality, particularly when employing open-source LLMs, we integrate Reinforcement Learning from Task Feedback (RLTF) (Ge et al., 2023a) following the application of our Formal-LLM framework. After the LLM-based agent generates a plan for a training task, the plan is executed on benchmark data to assess its performance. Then, the performance is used as a reward for RL to update the parameters of the LLM. Our framework ensures that invalid plans are always excluded during the LLM-based agent’s plan generation. Therefore, the number of valid rewards for LLM fine-tuning increases, and our approach can improve the performance of fine-tuned LLM-based agents.
@@ 5 Experiments
5.1 Backbone Large Language Model (LLM)
We test our Formal-LLM framework on these closed-source LLMs:
  • GPT-3.5-turbo (Brown et al., 2020) is a generative pre-trained transformer of OpenAI.
  • Claude-2 (Claude-2, 2023) is a transformer LLM of Anthropic.
  • GPT-4 (OpenAI, 2023) is a follow-up version of GPT-3.5.
and these open-source LLMs:
  • Flan-T5-Large (Chung et al., 2022) is a language model with 770 million parameters.
  • Vicuna-7B (Chiang et al., 2023) is a 7-billion size chatbot trained by fine-tuning the LLaMA model (Touvron et al., 2023a).
  • LLaMA-2-13B (Touvron et al., 2023b) is a successor to the 13-billion size LLaMA model.
@@ 5.2 Learning Schema of LLMs
  • Few-shot Learning (Few) presents a set of high-quality examples in the prompt, each consisting of input and desired output on the target task.
  • Reinforcement Learning from Task Feedback (RLTF) applies text constraint generation, executes the plan, and takes its performance as the reward to optimize the LLM by RL.
  • Formal-LLM (F-LLM) is our proposed framework, using automaton to control planning.
  • Formal-LLM plus RLTF (F-LLM+RLTF) applies our Formal-LLM framework on top of RLTF without text constraint generation. We use automaton to exclude invalid plans.
Specifically, we utilize Zero, Few, and F-LLM frameworks for closed-source LLMs, as these frameworks do not need to modify the LLM’s parameters. Regarding open-source LLMs, we compare RLTF, F-LLM and F-LLM+RLTF, given that RLTF surpasses Zero and Few, as exhaustively discussed in recent work (Ge et al., 2023a).
@@
5.3 Experimentation Datasets
5.3.1 Benchmark Datasets
We conduct experiments on two benchmark datasets, i.e., OpenAGI (Ge et al., 2023a) and TravelPlanner (Xie et al., 2024). Due to space limit, we provide the results on OpenAGI in the following and provide the results on TravelPlanner in Appendix Section A.4. The OpenAGI benchmark tasks are categorized based on their output type and ground-truth label type (Task 1, 2, and 3). Then, based on different task types, different metrics are employed to gauge the performance: CLIP Score (Hessel et al., 2021), assessing the similarity between text and image, is utilized for Text-to-Image tasks; BERT Score (Zhang et al., 2020), evaluating text generation with BERT, is applied when both data labels and the expected outputs are texts; and ViT Score (Wu et al., 2020) gauges the similarity between the image label and image output. Additionally, we construct Task X, which is a subset of “Task 1 $\cup$ Task 2 $\cup$ Task 3” that requires a tree-structured plan due to the use of many-input-single-ouput tools, such as Question Answering. Task X is used to test the complex planning ability of our Formal-LLM framework, and is evaluated using the corresponding metric. The tools are detailed in Table 3 and example tasks of each category are in Table 4 in the Appendix.
5.3.2 Real-life Practical Tasks
We also experiment with real-life planning scenarios, encompassing daily plans, cooking recipes, and risk management, where validity and rationality are crucial. In these scenarios, the concept of tool is generalized to include various types of steps in a plan, such as events, actions, or activities, as they can take diverse forms supporting the execution of the plan. For example, a tool in the daily planning task may be an activity such as taking breakfast. We provide qualitative analyses between the Zero and F-LLM learning schema using GPT-4 backbone for practical tasks. We use GPT-4 for the experiment of practical tasks because other LLMs can hardly generate readable plans for these tasks, and we test under Zero and F-LLM due to the limited sample size and that these two learning frameworks do not require accessing to the LLM parameters.
@@ 5.4 Experimental Analysis
The experiment results on the benchmark tasks are shown in Table 1 and Table 2, referring to the closed-source and open-source LLMs, respectively. Each row stands for a type of task, each column represents the learning schema of an LLM-based agent, and every three columns are the results of the same LLM. For closed-source LLMs, among the three learning schemata without revising the LLM parameter, almost all the best scores under each type of task belong to our F-LLM framework. For open-source LLMs, our F-LLM without fine-tuning is already better than RLTF (the best schema in the OpenAGI platform (Ge et al., 2023a)) in most cases, except for the ViT Score (Task 3), because the Type-3 tasks take the most significant portion of the RLTF’s fine-tuning data, and thus RLTF is adequately optimized on these tasks. The performance gain of our F-LLM framework comes from the 100% executable plans and supporting tree-structured planning for the difficult tasks in Task X. As a comparison, the best open-source LLM (LLaMA-2-13B with RLTF) can only generate 47% executable plans, and the best closed-source LLM (GPT-4 with few shot) can generate 76% executable plans. Additionally, none of the baselines can handle the Task X (Score = 0.0). Due to the increased amounts of valid rewards from the 100% executable plans with our framework, the F-LLM+RLTF approach enables boosted performance, as demonstrated by its better scores in Table 2. To summarize, the benchmark experiment results show that our Formal-LLM framework is an effective method to integrate the benefits of both natural and formal languages for more controllable and valid planning of LLM-based agents.
@@ 5.5 Case Study
We show the complete results of real-life examples by applying our Formal-LLM to the GPT models in the Appendix. From the results, after applying our Formal-LLM, the generated plan from the GPT-based agent is more complete, reasonable and specific to the case. In the daily plan example, the agent fails to fit all the activities into the plan without the Formal-LLM, while it can achieve this goal by applying the Formal-LLM. Still take Figure 4 as an example, our Formal-LLM can limit the planning to 10:00$\sim$20:00, while without Formal-LLM, the agent schedules activities after 20:00, even if we mention the constraint “Generate a plan for all the activities between 10:00 and 20:00” in natural language, which shows the advantages of formal language guided planning. For the cooking recipe example, without strict automaton constraint, GPT may generate steps such as “stir-fry the Chinese broccoli until it changes color” after “once the water boils, add the Chinese broccoli”. This is unreasonable because broccoli cannot be stir-fried in boiling water. For the risk management example, the generated plan without our framework is too general that it could be used for any other two companies. However, the plan after applying Formal-LLM is more specific and focuses on the potential antitrust risk between Microsoft and Blizzard. Thus, the Formal-LLM can use the automaton describing the constraints to generate higher-quality plans.
@@ 6 Conclusions and Future Work
In this study, we introduce the innovative Formal-LLM framework for LLM-based agents, combining LLM’s power of natural language comprehension with the precision of formal language. Our experiments, encompassing benchmark tasks and real-life practical scenarios, affirm the feasibility and effectiveness of employing automaton to control the agent’s generation of valid plans. More controllable LLM-based agents can augment the potential utilization of LLM in applications where high validity of planning is important.
Our work has several potential extensions. First, automating the translation of natural language into formal language for agents could further improve the framework.Additionally, this work focuses on LLM plan generation based on formal language. Another important problem to explore is LLM plan verification based on formal language.
Acknowledgment
We thank Harun Taha Kepenek for providing helpful suggestions to improve this research paper.
Appendix A Appendix
A.1 OpenAGI Benchmark Tasks and Tools
@@
Image Image ($A$) Colorization ($a_{1}$) ColorizerFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Zhang et al., 2017)
Super-Resolution ($a_{2}$) Swin2SRFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Conde et al., 2022)
Image Denoising ($a_{3}$) RestormerFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Zamir et al., 2022)
Image Deblurring ($a_{4}$) RestormerFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Zamir et al., 2022)
Text ($B$) Image Classification ($b_{1}$) ViTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Dosovitskiy et al., 2021)
Object Detection ($b_{2}$) DETRFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Carion et al., 2020)
Image Captioning ($b_{3}$) Vision Encoder DecoderFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Li et al., 2023b)
Text Image ($C$) Text-to-Image Generation ($c_{1}$) Stable DiffusionFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Rombach et al., 2022)
Text ($D$) Sentiment Analysis ($d_{1}$) DistilBERTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Sanh et al., 2020)
Text Summarization ($d_{2}$) BARTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Lewis et al., 2019)
Machine Translation ($d_{3}$) T5Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Raffel et al., 2020)
Fill Mask ($d_{4}$) DistilBERTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Sanh et al., 2020)
Text Generation ($d_{5}$) GPT-2Formal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Radford et al., 2019)
Image-Text Pair Text ($E$) Visual Question Answering ($e_{1}$) GITFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Wang et al., 2022)
Text-Text Pair Text ($F$) Question Answering ($f_{1}$) DistilBERTFormal-LLM: Integrating Formal Language and Natural Language for Controllable LLM-based Agents (Sanh et al., 2020)
@@
A.2 Implementation Details
Our framework and all baselines are implemented by PyTorch, an open-source library. We follow the implementation setting of the OpenAGI platform (Ge et al., 2023a) for baselines. For the F-LLM+RLTF framework, we use the REINFORCE (Williams, 1992) as the core reinforcement learning algorithm of RLTF. We use the original checkpoint of each backbone LLM without supervised fine-tuning. We set the maximum number of updating epochs at 30 and use Adam as the optimizer with the learning rate at 0.001 for RLTF. Also, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to the RLTF for efficient fine-tuning with the rank as 8 and the placement as $v_{proj}$ ($W_{v}$) and $q_{proj}$ ($W_{q}$). Our experimental results are the average of five runs.
A.3 Prompt for OpenAGI Benchmark Tasks
An example of the CFG and the corresponding PDA are shown in Eq.(3) and Eq.(4) as well as Figure 3. Different benchmark tasks share a common subset of constraints shown in Eq.(3) corresponding to the input-output constraints of the tools. Besides, each task has its own constraint on the input-output of the task, resulting in different constraints in Eq.(4), and thus the final PDA (Figure 3) for different tasks are different. Also, we directly use the tool name (e.g., Colorization) instead of the category name (e.g., “Image-in, Image-out tools”) in the prompt. For example, consider the PDA in Figure 3 when it is in state $q_{1}$ with the symbol $I$ at the top of the stack. There are three viable transits: $(\varepsilon,I;AI)$, $(\varepsilon,I;CT)$, and $(i,I;\varepsilon)$. In practice, as nonterminals $A$ and $C$ are promptly replaced by $terminals$ from $a_{1}\sim a_{4}$ and $c_{1}$. This is because the tool name provides more concrete information for the LLM to comprehend the tool’s functionality, enabling the LLM to make informed decisions without influencing the plan’s executability.